# Approximate Nearest Neighbor Search with Window Filters

### v1

@@ Abstract
We define and investigate the problem of c-approximate window search: approximate nearest neighbor search where each point in the dataset has a numeric label, and the goal is to find nearest neighbors to queries within arbitrary label ranges. Many semantic search problems, such as image and document search with timestamp filters, or product search with cost filters, are natural examples of this problem. We propose and theoretically analyze a modular tree-based framework for transforming an index that solves the traditional c-approximate nearest neighbor problem into a data structure that solves window search. On standard nearest neighbor benchmark datasets equipped with random label values, adversarially constructed embeddings, and image search embeddings with real timestamps, we obtain up to a 75×75$\times$75 × speedup over existing solutions at the same level of recall.
@@ 1 Introduction
Recent advances in vector embeddings for text, images, and other unstructured data have led to a surge in research and industry interest in ANNS. This trend is reflected by a proliferation of commercial vector databases: Pinecone (Pinecone Systems, Inc., 2024), Vearch (vea, 2022), Weaviate (wea, 2022), Milvus (mil, 2022), and Vespa (ves, 2022), among others. These databases offer hyperparameter tuning, metadata storage, and solutions to a variety of related search problems. They especially tout their support of ANNS with metadata filtering as a key differentiating application.
In this work, we are primarily interested in a difficult generalization of the $c$-approximate nearest neighbor problem that we term Window Search (see Definition 3.3). Window search is similar to ANNS, except queries are accompanied by a “window filter”, and the goal is to find the nearest neighbor to the query that has a label value within the filter. This problem has a large number of immediate applications. For instance, in document and image search, each item may be accompanied by a timestamp, and the user may wish to filter to an arbitrary time range (e.g. they may wish to search for hiking pictures, but only from last summer, or forum posts about a bug, but only in the days after a new version was released). Another application is product catalogs, where a user may wish to filter search results by cost. Finally, we note that a large class of emerging applications is large language model retrieval augmented generation, where by storing and retrieving information LLMs can reduce hallucinations and improve their reasoning (Peng et al., 2023); window search may be critical when the LLM needs to recall something stored on a certain date or range of dates.
@@
    We prove our tree-based framework solves window search and give runtime bounds for the general case and for a specific instantiation with DiskANN (Jayaram Subramanya et al., 2019). We also analyze optimal partitioning strategies for the label-space partitioning approach.
@@
    We benchmark our methods against strong baselines and vector databases, achieving up to a 75×75$\times$75 × speedup on real world embeddings and adversarially constructed data.
@@ 2 Related Work
Another popular direction focuses on dedicated indices for filtered ANNS, which consistently outperform their general-purpose counterparts (Simhadri et al., 2023). For example, FilteredDiskANN, an adaptation of DiskANN (Jayaram Subramanya et al., 2019) that supports filters that are conjunctive ORs of one or more boolean labels, builds a graph that can be traversed by a modified beam search excluding points not matching the filter (Gollapudi et al., 2023). Another approach is CAPS, which is an inverted index where each partition stores a Huffman tree dividing points by their labels (Gupta et al., 2023). While these methods are the state of the art, they are restricted to filtering on boolean labels; they do not support window search.
Segment Trees. Segment trees (and the closely related Fenwick tree (Fenwick, 1994)) are tree data structures built over an array that recursively sub-divide the array to obtain a balanced binary tree (Bentley, 1977). By storing appropriate augmented values at the internal nodes of this tree, these data structures can be used to support a variety of queries over arbitrary intervals in the array, e.g., computing the maximum value in any given query interval $[l,r]$. Segment and Fenwick trees can be generalized to higher fanout trees, i.e., $B$-ary segment or Fenwick trees that have a fanout of $B$ and a height of $\lceil\log_{B}n\rceil$ (Pibiri & Venturini, 2021). Our work adapts these tree structures to the window search problem by designing a similar data structure that stores ANNS indices at internal tree nodes. Huang et al. (2023) used the Fenwick tree for filtered search within a clustering context, but their work only considers a prefix interval ($[0,r]$), and they used $k$d-trees, which are designed for low-dimensional data.
Filtered Search in Relational Databases. Traditional relational database systems support arbitrary range-based queries by constructing B-trees or log-structured merge trees (Ilyas et al., 2008; Qader et al., 2018). These databases have complex query planning systems that determine during execution whether and how to query these constructed indices (see, e.g., (Kurc et al., 1999)), but do not typically have support for ANNS.
@@
$A$ Arbitrary $c$-ANN algorithm, e.g., DiskANN
@@
$\alpha$ Pruning parameter for DiskANN
@@ 4 Window Search Algorithms
In this section, we describe algorithms to solve the window search problem. In Section 4.1, we examine two naive baselines for solving window search; in Section 4.2, we introduce a new data structure, the $\beta$-WST, and design an algorithm to query it; and in Section 4.3, we examine additional algorithms for solving window search.
We note that with a fixed window filter $(a,b)$, a reasonable approach to solving window search is simply to index $D_{(a,b)}$ using an off-the-shelf ANNS algorithm and then query it with each $q$. Thus, we are interested in the more challenging problem where we receive queries with arbitrary window filters.
@@ 4.1 Naive Baselines
$\mathsf{Prefiltering}$ is a naive baseline that works by sorting all of the points by label ahead of time. Given a query $x$ with window filter $(a,b)$, we do binary search on the sorted labels to find the start and end of the range that meet the filter constraints, and then find the distance between $x$ and every point in the range and return the closest point.
$\mathsf{Postfiltering}$ (Yu et al., 2023; Chen et al., 2020) is a second baseline that works by first building an index on all of $D$ using an ANN algorithm $A$. To do a window search, we query $A$ for $k=1$ point, and then continually double $k$ until at least one point is returned that has a label within $(a,b)$, and then we return the closest of these points. We additionally define a hyperparameter called $\mathsf{final\_multiply}$; if this is greater than $1$, then we perform a final additional search with a $\mathsf{final\_multiply}$ times larger value of $k$.
Figure 1: Top left is a $2$-WST. Each index contains a recursive partition of the entire dataset $D$, indexed into a graph for fast ANN (see Algorithm 1). Top right shows the structure of an example label space partitioning method that ensures no optimized postfiltering query will have a large blowup (see Theorem 5.6). Bottom shows different query methods; from left to right: a tree-based query as in Algorithm 2, an optimized postfiltering query with a small blowup (see Definition 5.4), and an optimized postfiltering query with a large blowup (see Definition 5.4).
@@ 4.2 $\beta$-Window Search Tree
Index Construction. A $\beta$-WST $T$ can be constructed as shown in Algorithm 1. In the base case, if the dataset $D$ is small, we do not construct any tree node (Lines 34). Otherwise, we construct an ANN index of $D$ (Line 5). On lines 67 we define the sizes for splitting $D$ into $\beta$ partitions, all of which are equally sized except the last, which may be smaller. On line 8 we initialize the children nodes. We then loop through every partition in parallel (Lines 912) and recursively call BuildTree on the set of points (sorted by label) corresponding to the start and end of the partition. Finally, on line 13 we return the constructed tree, which is a tuple consisting of an ANN index built on $D$, the result of BuildTree called on each child with the size of each child, and the point set $D$.
@@
5:  $\texttt{index}\leftarrow A(D)$
@@
Querying the Index. We query a $\beta$-WST using Algorithm 2. The input is a $\beta$-WST $T$ as built by Algorithm 1, a query point $q$ , and a window filter $(a,b)$. At a high level, Algorithm 2 recurses through the tree constructed by Algorithm 1 and queries instances of $A$ that union together to equal the entire filtered dataset. If the dataset is small and the index is a leaf node NULL, we do a brute force search over $D$ (Lines 34). If the window filter $(a,b)$ covers all points, we query the ANN index and return the result (Lines 56). Otherwise, $D$ has some points that do not have a label in $(a,b)$, so we loop over each child (Line 8) and recurse into it if some of the points in the child meet the query’s window filter constraint. Finally, for each one of these children that we recurse into we add the returned points to a candidate list $\mathcal{L}_{\text{cand}}$, and return the closest point from $\mathcal{L}_{\text{cand}}$ on line 13.
@@ 4.3 Additional Query Methods
In addition to Algorithm 2 above, we examine a number of additional query methods for window search that come with various trade offs.
  • $\mathsf{OptimizedPostfiltering}$ uses the index built by Algorithm 1 but uses a novel query algorithm. Given a query $x$ with window filter $(a,b)$, $\mathsf{OptimizedPostfiltering}$ finds the smallest subset $S$ of $D$ corresponding to an index we built $I=A(S)$ where $D_{(a,b)}\subset S$, and then queries that index using the same procedure as described in $\mathsf{Postfiltering}$. A “small blowup” query is one in which the smallest subset $S$ is not that much larger than $D_{(a,b)}$, whereas a “large blowup” query is one where $S$ is much larger than $D_{(a,b)}$. These small and large blowup queries are shown on the bottom right of Figure 1. Blowup factor is defined formally in Definition 5.4.
  • $\mathsf{ThreeSplit}$ also uses the index built by Algorithm 1 and a novel querying algorithm. Similar to Algorithm 2, a query initially finds the highest level where any partition at all is entirely contained in the window filter, and then does a query on every one of these partitions. Instead of recursing further down the tree, however, $\mathsf{ThreeSplit}$ then does an $\mathsf{OptimizedPostfiltering}$ call on each of the remaining label subranges on each side of the middle “covered” label portion. Because we fill in the middle first, we are guaranteed that there can be no ”large blowup” case like in $\mathsf{OptimizedPostfiltering}$.
@@ 5 Theoretical Analysis
5.1 Analysis of Querying a $\beta$-WST (Algorithm 2)
In our analysis in this section, we assume without loss of generality that $N$ is a power of $\beta$. Removing this assumption would lead to more floor and ceiling operators in Theorem 5.1 and leave our other results unchanged. In the interest of space, we defer proofs to the Appendix.
@@ 5.1 Analysis of Querying a $\beta$-WST (Algorithm 2)
Many theoretical ANN results in the literature have a runtime of $O(N^{\rho})$ for constant $\rho$, e.g., LSH (Andoni & Indyk, 2008) and $k$-nearest neighbor graphs (Prokhorenkova & Shekhovtsov, 2020). Other results have a runtime that is paramterized only with constants describing the data distribution, and have no reliance on $N$. By applying Theorem 5.1, we have the following results for these common function classes:
@@
Finally, we can apply Lemma 5.2 to a recent $c$-ANN result for DiskANN graph-based search (Indyk & Xu, 2023). This is particularly relevant because we use DiskANN graphs for our experiments. Letting $\Delta$ be the aspect ratio of $D$, i.e., the ratio between the maximum and minimum distances between any two pairs of points, we have the following result:
@@
Algorithm 2 instantiated with a “slow preprocessed” $\alpha$-DiskANN graph solves the $c$-approximate window search problem in any metric space on a dataset with doubling dimension $\delta$ and aspect ratio $\Delta$ in running time

### v2

@@ Abstract
We define and investigate the problem of c-approximate window search: approximate nearest neighbor search where each point in the dataset has a numeric label, and the goal is to find nearest neighbors to queries within arbitrary label ranges. Many semantic search problems, such as image and document search with timestamp filters, or product search with cost filters, are natural examples of this problem. We propose and theoretically analyze a modular tree-based framework for transforming an index that solves the traditional c-approximate nearest neighbor problem into a data structure that solves window search. On standard nearest neighbor benchmark datasets with random label values, adversarially constructed embeddings, and image search embeddings with real timestamps, we obtain up to a 75×75$\times$75 × speedup over existing solutions at the same level of recall.
@@ 1 Introduction
Recent advances in vector embeddings for text, images, and other unstructured data have led to a surge in research and industry interest in ANNS. This trend is reflected by a proliferation of commercial vector databases: Pinecone (Pinecone Systems, Inc., 2024), Vearch (vea, 2022), Weaviate (wea, 2022), Milvus (mil, 2022), and Vespa (ves, 2022), among others. These databases offer hyper-parameter tuning, metadata storage, and solutions to a variety of related search problems. They especially tout their support of ANNS with metadata filtering as a key differentiating application.
In this work, we are primarily interested in a difficult generalization of the $c$-approximate nearest neighbor problem that we term Window Search (see Definition 3.3). Window search is similar to ANNS, except queries are accompanied by a “window filter”, and the goal is to find the nearest neighbor to the query that has a label value within the filter. This problem has a large number of immediate applications. For instance, in document and image search, each item may be accompanied by a timestamp, and the user may wish to filter to an arbitrary time range (e.g., they may wish to search for hiking pictures, but only from last summer, or forum posts about a bug, but only in the days after a new version was released). Another application is product catalogs, where a user may wish to filter search results by cost. Finally, we note that a large class of emerging applications is large language model retrieval augmented generation, where by storing and retrieving information LLMs can reduce hallucinations and improve their reasoning (Peng et al., 2023); window search may be critical when the LLM needs to recall something stored on a certain date or range of dates.
@@
    We prove that our tree-based framework solves window search and give runtime bounds for the general case and for a specific instantiation with the Vamana ANNS algorithm (Jayaram Subramanya et al., 2019). We also analyze optimal partitioning strategies for the label-space partitioning approach.
@@
    We benchmark our methods against strong baselines and vector databases, achieving up to a 75×75$\times$75 × speedup on real world embeddings and adversarially constructed datasets.
@@ 2 Related Work
Another popular direction focuses on dedicated indices for filtered ANNS, which consistently outperform their general-purpose counterparts (Simhadri et al., 2023). For example, FilteredDiskANN, an adaptation of DiskANN (Jayaram Subramanya et al., 2019) that supports filters that are conjunctive ORs of one or more boolean labels, builds a graph that can be traversed by a modified beam search excluding points not matching the filter (Gollapudi et al., 2023). Another approach is CAPS, which is an inverted index where each partition stores a Huffman tree dividing points by their labels (Gupta et al., 2023). While these methods are the state of the art, they are restricted to filtering on boolean labels; they do not support window search. Finally, recent work (Mohoney et al., 2023) presents a tree based nearest neighbor engine for combined vector-metadata searches and show range-based filters as an application, but their partitioning requires historical queries and many of their gains come from batching queries to avoid redundant computation. Their method also only builds ANNS indices in the “leaves” of the tree, whereas we build indices at internal nodes, which is the key idea that ensures our method only needs to query a logarithmic number of ANNS indices. Thus, although their code is not open source, we do not expect their system to perform as well as ours on window search.
Segment Trees. Segment trees (and the closely related Fenwick trees (Fenwick, 1994)) are tree data structures built over an array that recursively sub-divide the array to obtain a balanced binary tree (Bentley, 1977). By storing appropriate augmented values at the internal nodes of this tree, these data structures can be used to support a variety of queries over arbitrary intervals in the array, e.g., computing the maximum value in any given query interval $[l,r]$. Segment and Fenwick trees can be generalized to higher fanout trees, i.e., $B$-ary segment or Fenwick trees that have a fanout of $B$ and a height of $\lceil\log_{B}n\rceil$ (Pibiri & Venturini, 2021). Our work adapts these tree structures to the window search problem by designing a similar data structure that stores ANNS indices at internal tree nodes. Huang et al. (2023) use the Fenwick tree for filtered search within a clustering context, but their work only considers a prefix interval ($[0,r]$), and they use $k$d-trees, which are designed for low-dimensional data.
Filtered Search in Relational Databases. Traditional relational database systems support arbitrary range-based queries by constructing B-trees or log-structured merge trees (Ilyas et al., 2008; Qader et al., 2018). These databases have complex query planning systems that determine during execution whether and how to query these constructed indices (see, e.g., (Kurc et al., 1999)), but typically do not have support for ANNS. The few that do have support for ANNS use an existing open-source ANNS implementation to perform the search (e.g., pgvector (Kane, 2024), an ANNS add-on for PostgreSQL, uses HNSW (Malkov & Yashunin, 2018).
Table 1: Notation used in this paper.
@@
$A$ Arbitrary $c$-ANN algorithm, e.g., Vamana
@@
$\alpha$ Pruning parameter for Vamana
@@ 4 Window Search Algorithms
In this section, we describe algorithms for solving the window search problem. In Section 4.1, we examine two naive baselines for solving window search; in Section 4.2, we introduce a new data structure, the $\beta$-WST, and design an algorithm to query it; and in Section 4.3, we examine additional algorithms for solving window search.
We note that with a fixed window filter $(a,b)$, a reasonable approach to solving window search is simply to index $D_{(a,b)}$ using an off-the-shelf ANNS algorithm and then query it with each $q$. Thus, we are interested in the more challenging problem where queries have arbitrary window filters.
@@ 4.1 Naive Baselines
As we discuss in Section 2, prefiltering and postfiltering are the current state of the art for filtered search. Here, we describe the specific way we implement them in more detail, as they are the main baselines that we compare against in our experiments.
$\mathsf{Prefiltering}$ is a naive baseline that works by sorting all of the points by label ahead of time. Given a query $x$ with window filter $(a,b)$, we perform binary search on the sorted labels to find the start and end of the range that meet the filter constraints, and then find the distance between $x$ and every point in the range and return the closest point.
$\mathsf{Postfiltering}$ (Yu et al., 2023; Chen et al., 2020) is a second baseline that works by first building an index on all of $D$ using an ANNS algorithm $A$. To perform a window search, we query $A$ for $k\geq 1$ points, and then repeatedly double $k$ until at least one point is returned that has a label within $(a,b)$. Finally, we return the closest of these points. We additionally define a hyper-parameter called $\mathsf{final\_multiply}$; if this is greater than $1$, then we perform a final additional search with a $\mathsf{final\_multiply}$ times larger value of $k$.
Figure 1: Top left: A $2$-WST. Each node of the tree in green contains a recursive partition of the entire dataset $D$ indexed by an ANNS algorithm (see Algorithm 1). The graph in each green node represents a graph-based ANNS index built by an algorithm like Vamana. Top right: the structure of an example label space partitioning method that ensures that no optimized postfiltering query will have a large blowup (see Theorem 5.7). Bottom: Different query methods; from left to right: a tree-based query as in Algorithm 2, an optimized postfiltering query with a small blowup (see Definition 5.5), and an optimized postfiltering query with a large blowup (see Definition 5.5).
@@ 4.2 $\beta$-Window Search Tree
Index Construction. A $\beta$-WST $T$ can be constructed as shown in Algorithm 1. In the base case, if the dataset $D$ is small, we do not construct any tree node (Lines 34). Otherwise, we construct an ANNS index of $D$ (Line 5). On Lines 67, we define the sizes for splitting $D$ into $\beta$ partitions, all of which are equally sized except the last, which may be smaller. On Line 8, we initialize the children nodes. We then loop through every partition in parallel (Lines 912) and recursively call BuildTree on the set of points (sorted by label) corresponding to the start and end of the partition. Finally, on Line 13 we return the constructed tree, which is a tuple consisting of an ANNS index built on $D$, the result of BuildTree called on each child along with the size of each child, and the point set $D$.
@@
5:  $\texttt{index}\leftarrow A(D)$The next two lines define the sizes of the $\beta$ subsets; all are equal size except the last}
@@
Querying the Index. We query a $\beta$-WST using Algorithm 2. The input is a $\beta$-WST $T$ as built by Algorithm 1, a query point $q$ , and a window filter $(a,b)$. At a high level, Algorithm 2 recurses through the tree constructed by Algorithm 1 and queries instances of $A$ that union together to equal the entire filtered dataset. If the dataset is small and the index is a leaf node NULL, we do a brute force search over $D$ (Lines 34). If the window filter $(a,b)$ covers all points, we query the ANNS index and return the result (Lines 56). Otherwise, $D$ has some points that do not have a label in $(a,b)$, so we loop over each child (Line 8) and recurse into it if some of the points in the child meet the query’s window filter constraint. Finally, for each one of these children that we recurse into, we add the returned points to a candidate list $\mathcal{L}_{\text{cand}}$, and return the closest point from $\mathcal{L}_{\text{cand}}$ on Line 13.
@@ 4.3 Additional Query Methods
In addition to Algorithm 2 above, we examine a number of additional query methods for window search that come with various trade-offs.
  • $\mathsf{OptimizedPostfiltering}$ uses the index built by Algorithm 1 but uses a novel query algorithm, shown in Algorithm 3. Given a query $x$ with window filter $(a,b)$, $\mathsf{OptimizedPostfiltering}$ finds the smallest subset $S$ of $D$ corresponding to an index that we built $I=A(S)$ where $D_{(a,b)}\subset S$, and then queries that index using the same procedure as described in $\mathsf{Postfiltering}$. A “small blowup” query is one in which the smallest subset $S$ is not that much larger than $D_{(a,b)}$, whereas a “large blowup” query is one where $S$ is much larger than $D_{(a,b)}$. These small and large blowup queries are shown on the bottom right of Figure 1. Blowup factor is defined formally in Definition 5.5.
  • $\mathsf{ThreeSplit}$ also uses the index built by Algorithm 1 and a novel querying algorithm, shown in Algorithm 4. Similar to Algorithm 2, a query initially finds the highest level where any partition at all is entirely contained in the window filter, and then does a query on every one of these partitions. Instead of recursing further down the tree, however, $\mathsf{ThreeSplit}$ then does an $\mathsf{OptimizedPostfiltering}$ call on each of the remaining label subranges on each side of the middle “covered” label portion. Because we fill in the middle first, we are guaranteed that there can be no ”large blowup” case like in $\mathsf{OptimizedPostfiltering}$.
@@
Algorithm 3 OptimizedPostfiltering$(T,q,(a,b))$
1:  Input: $\beta$-WST $T=(\texttt{index},(\texttt{children},\texttt{sizes}),D)$, query $q$, window filter $(a,b)$.
[…]
4:  $k\leftarrow 1$
5:  while $k<N$ do
6:    $\mathcal{L}_{\text{unfiltered}}\leftarrow\texttt{index}(q,k)$
7:    $\mathcal{L}_{\text{cand}}\leftarrow\{x\in\mathcal{L}_{\text{cand}}\;|\>\ell(x)\in(a,b)\}$
8:    if $\mathcal{L}_{\text{cand}}\neq\emptyset$ then
9:       return $\operatorname*{arg\,min}_{y\in\mathcal{L}_{\text{cand}}\cap D_{(a,b)}}\texttt{dist}_{V}(q,y)$
10:    $k\leftarrow 2k$
11:  return $\emptyset$
Algorithm 4 ThreeSplit$(T,q,(a,b))$
1:  Input: $\beta$-WST $T=(\texttt{index},(\texttt{children},\texttt{sizes}),D)$, query $q$, window filter $(a,b)$.
3:  $\texttt{index},(a^{\prime},b^{\prime})\leftarrow$ index in $T$ containing the most points with labels in $(a,b)$, label range of points in index
4:  $\mathcal{L}_{\text{cand}}=\text{OptimizedPostfiltering}(T,q,(a,a^{\prime}))\cup\texttt{index}(q)\cup\text{OptimizedPostfiltering}(T,q,(b^{\prime},b))$
5:  return $\operatorname*{arg\,min}_{y\in\mathcal{L}_{\text{cand}}\cap D_{(a,b)}}\texttt{dist}_{V}(q,y)$
@@ 5 Theoretical Analysis
5.1 Analysis of Building and Querying a $\beta$-WST
In our analysis in this section, we assume without loss of generality that $N$ is a power of $\beta$. Removing this assumption would lead to more floor and ceiling operators in Theorem 5.2 and leave our other results unchanged. We defer proofs to Appendix A. We will first analyze the construction time and memory of Algorithm 1, and then we will prove the correctness and analyze the query time of Algorithm 2.
Consider some function $A_{f}$ parameterized by the dataset $D$, dataset size $N$, and subset size $m$. For example, $A_{f}$ may be a construction time function or a memory function. This function evaluated on all nodes of a $\beta$-WST is
$O\left(\sum_{j=0}^{\log_{\beta}N}\beta^{j}A_{f}(D,N\cdot\beta^{-j})\right).$
Since we use Vamana indices in our experiments, we now apply this result to obtain the running time of Algorithm 1 (the construction time) and the memory of the resulting index. We use a recent analysis for Vamana graph-based search (Indyk & Xu, 2023) that assumes a “slow-preprocessing” index construction with runtime $O(N^{3})$ and memory $O(N(4\alpha)^{\delta}\log\Delta)$. Letting $\Delta$ be the aspect ratio of $D$, i.e., the ratio between the maximum and minimum distances between any two pairs of points, we have the following result:
Lemma 5.1.
Algorithm 1 instantiated with a “slow preprocessed” $\alpha$-Vamana graph runs in time
$O\left(\frac{1}{1-\beta^{-2}}N^{3}\right)=O(N^{3})$
and returns a $\beta$-WST of memory
$O\left((4\alpha)^{\delta}\log(\Delta)N\log_{\beta}N\right).$
We now move on to our main runtime theorem, which both proves that Algorithm 2 indeed solves $c$-approximate window search and upper bounds the running time for an arbitrary ANNS index $A$:
@@ 5.1 Analysis of Building and Querying a $\beta$-WST
Many theoretical ANN results in the literature have a runtime of $O(N^{\rho})$ for a constant $\rho$, e.g., LSH (Andoni & Indyk, 2008) and $k$-nearest neighbor graphs (Prokhorenkova & Shekhovtsov, 2020). Other results have a runtime that is parameterized only with constants describing the data distribution, and have no reliance on $N$. By applying Theorem 5.2, we have the following results for these common function classes:
@@
We can again apply Lemma 5.3 using recent $c$-ANN results for Vamana graph-based search (Indyk & Xu, 2023). We have the following result:
@@
Algorithm 2 instantiated with a “slow preprocessed” $\alpha$-Vamana graph solves the $c$-approximate window search problem in any metric space on a dataset with doubling dimension $\delta$ and aspect ratio $\Delta$ in running time